(2 points) Write the equation for the line y - 6 = 5 (x - 7) in the form y = m x + b, and enter your answer in this form.

Answer by mananth(16946)

y - 6 = 5 (x - 7)
y-6 =5x-35
y= 5x -35+6
y= 5x-29